OpenStudy (ccastudentgrade10):

FAN AND MEDAL find a quadratic function in standard form for each set of points (0,-4), (1,0), (2,2)

OpenStudy (ccastudentgrade10):

@xMissAlyCatx @texaschic101 @dayakar @zepdrix

OpenStudy (xmissalycatx):

Do you know what quadratic function and standard form are?

OpenStudy (ccastudentgrade10):

yes, i learned them a long time ago but idk how to make these points like them

OpenStudy (ccastudentgrade10):

im v confused @xMissAlyCatx

OpenStudy (ccastudentgrade10):

how do i get a b and c?

OpenStudy (xmissalycatx):

@Compassionate @ParthKohli I honestly don't know how to explain this so here you go..

OpenStudy (ccastudentgrade10):

anyone? i really need help with this @texaschic101 @dayakar @zepdrix

OpenStudy (dayakar):

\[y = ax ^{2}+bx +c\]

OpenStudy (ccastudentgrade10):

okay but i still dont understand how you get a, b, and c

OpenStudy (dayakar):

(x1,y1) =(0, -4), (x2, y2) =(1,0) , (x3,y3) = (2,2) plug each one in the above equation u can get 3 equations

OpenStudy (dayakar):

-4 = a* 0^2 +b*0 +c[ plug in first one] -4 =c c= -4 -------(1)

OpenStudy (dayakar):

do u understand

OpenStudy (dayakar):

plz , respond

OpenStudy (dayakar):

0 = a*1^2 +b* 1 +c [ plug in second in the equation] 0 = a+ b -4 [ from (1) c= -4] ------(2)

OpenStudy (ccastudentgrade10):

sorry my wifi was glitching

OpenStudy (dayakar):

2= a*2^2 + b*2 +c [ plug in third] 2 = 4a +2b -4 [ from (1) ,c= -4] divide each term with 2 u can get 1= 2a +b -2 -----(3) subtract (2) from (3) u can get a value

OpenStudy (ccastudentgrade10):

1?

OpenStudy (ccastudentgrade10):

@dayakar

OpenStudy (ccastudentgrade10):

so does b = 1

OpenStudy (dayakar):

a= -1

OpenStudy (dayakar):

plug a=-1 in (2)

OpenStudy (dayakar):

is it clear

OpenStudy (ccastudentgrade10):

im confused where im plugging in a =-1

OpenStudy (ccastudentgrade10):

c = -4 right

OpenStudy (dayakar):

equation (2) 0= a+b-4

OpenStudy (ccastudentgrade10):

0=b-5?

OpenStudy (dayakar):

we get c = -4 very beginning of the solution

OpenStudy (ccastudentgrade10):

okay, so does b = 5?

OpenStudy (dayakar):

0 = -1 +b -4 0= -5+b 5=b b=5

OpenStudy (dayakar):

there fore a= -1, b= 5, c=-4 write in the standard form y= ax^2+bx+c y= -1x^2 +5x-4

OpenStudy (ccastudentgrade10):

OHHHH okay that makes sense

OpenStudy (ccastudentgrade10):

if i post another could u help me?

OpenStudy (dayakar):

i'll try